# Trusted Multi-view Learning with Label Noise

### v1

@@ Abstract
Multi-view learning methods often focus on improving decision accuracy while neglecting the decision uncertainty, which significantly restricts their applications in safety-critical applications. To address this issue, researchers propose trusted multi-view methods that learn the class distribution for each instance, enabling the estimation of classification probabilities and uncertainty. However, these methods heavily rely on high-quality ground-truth labels. This motivates us to delve into a new generalized trusted multi-view learning problem: how to develop a reliable multi-view learning model under the guidance of noisy labels? We propose a trusted multi-view noise refining method to solve this problem. We first construct view-opinions using evidential deep neural networks, which consist of belief mass vectors and uncertainty estimates. Subsequently, we design view-specific noise correlation matrices that transform the original opinions into noisy opinions aligned with the noisy labels. Considering label noises originating from low-quality data features and easily-confused classes, we ensure that the diagonal elements of these matrices are inversely proportional to the uncertainty, while incorporating class relations into the off-diagonal elements. Finally, we aggregate the noisy opinions and employ a generalized maximum likelihood loss on the aggregated opinion for model training, guided by the noisy labels. We empirically compare TMNR with state-of-the-art trusted multi-view learning and label noise learning baselines on 5 publicly available datasets. Experiment results show that TMNR outperforms baseline methods on accuracy, reliability and robustness. We promise to release the code and all datasets on Github and show the link here.
@@ 1 Introduction
Multi-view data is widely present in various real-world scenarios. For instance, in the field of healthcare, a patient’s comprehensive condition can be reflected through multiple types of examinations; social media applications often include multi-modal contents such as textual and visual reviews. Multi-view learning synthesizes both consistency and complementary information to obtain a more comprehensive understanding of the data. It has generated significant and wide-ranging influence across multiple research areas, including classification Liang et al. (2021), clustering Xu et al. (2023); Huang et al. (2022), recommendation systems Lin et al. (2023); Nikzad-Khasmakhia et al. (2020), and large language models Min et al. (2023).
@@
Figure 1: Visualization of the generalized trusted multi-view learning problem: the model should recognize the feature and model uncertainty caused by low-quality feature and noisy label, respectively.
@@
Most existing multi-view learning methods focus on improving decision accuracy while neglecting the decision uncertainty. This significantly limits the application of multi-view learning in safety-critical scenes, such as healthcare. Recently, Han et al. propose a pioneering work Han et al. (2020), Trusted Multi-view Classification (TMC), to solve this problem. TMC calculates and aggregates the evidences of all views from the original data features. It then utilizes these evidences to parameterize the class distribution, which could be used to estimate the class probabilities and uncertainty. To train the entire model, TMC requires the estimated class probabilities to be consistent with the ground-truth labels. Following this line, researchers propose novel evidence aggregation methods, aiming to enhance the reliability and robustness in the presence of feature noise Gan et al. (2021); Qin et al. (2022), conflictive views Xu et al. (2024) and incomplete views Xie et al. (2023) et al.
@@
Regretfully, these trusted multi-view learning methods consistently rely on high-quality ground-truth labels. The labeling task is time-consuming and expensive especially when dealing with large scale datasets, such as user generated multi-modal contents in social media applications. This motivates us to delve into a new Generalized Trusted Multi-view Learning (GTML) problem: how to develop a reliable multi-view learning model under the guidance of noisy labels? This problem encompasses two key objectives: 1) detecting and refining the noisy label during the training stage; 2) recognizing the model’s uncertainty caused by noisy labels. For example, instances belonging to classes like “cat” and “wolf” might exhibit similarities and are prone to being mislabelled. Consequently, the model should exhibit higher decision uncertainty in such cases. An intuitive analogy is an intern animal researcher (model) may not make high-confidence decisions for all animals (instances), but is aware of the cases where a definitive decision is challenging.
@@
In this paper, we propose an Trusted Multi-view Noise Refining (TMNR) method for the GTML problem. We consider the label noises arising from two sources: low-quality data features, such as blurred and incomplete features, and easily-confused classes, such as classes “cat” and “wolf”. Our objective is to leverage multi-view consistent information for noise detection. To achieve this, we first construct the view-specific evidential Deep Neural Networks (DNNs) to learn view-specific evidence, which could be termed as the amount of support to each category collected from data. We then model the view-specific distributions of class probabilities using the Dirichlet distribution, parameterized with view-specific evidence. These distributions allow us to construct opinions, which consist of belief mass vectors and uncertainty estimates. We design view-specific noise correlation matrices to transform the original opinions to the noisy opinions, which aligns with the noisy label. Considering that low-quality data features are prone to mislabeling, we require the diagonal elements of the noise correlation matrices to be inversely proportional to the uncertainty. Additionally, we incorporate class relations into the off-diagonal elements. For instance, the elements corresponding to “cat” and “wolf” should have larger values since these two classes are easily mislabelled. Next, we aggregate the noisy opinions to obtain the common evidence. Finally, we employ a generalized maximum likelihood loss on the common evidence, guided by the noisy labels, for model training.
@@ 2.1 Deep Multi-view Fusion
Multi-view fusion has demonstrated its superior performance in various tasks by effectively combining information from multiple sources or modalities Xu et al. (2022); Zhou et al. (2023). According to the fusion strategy, existing deep multi-view fusion methods can be roughly classified into feature fusion Wang et al. (2015); Xu et al. (2022); Lin et al. (2022) and decision fusion Jillani et al. (2020); Liu et al. (2022). A major challenge of feature fusion methods is each view might exhibit different types and different levels of noise at different points in time. Trust decision fusion methods solve this making view-specific trust decisions to obtain the view-specific reliabilities, then assigning large weights to these views with high reliability in the multi-view fusion stage. Following this line, Xie et al. Xie et al. (2023) tackle the challenge of incomplete multi-view classification through a two-stage approach, involving completion and evidential fusion. Xu et al. Xu et al. (2024) focus on making trust decisions for instances that exhibit conflicting information across multiple views. They propose an effective strategy for aggregating conflicting opinions and theoretically prove this strategy can exactly model the relation of multi-view common and view-specific reliabilities. However, it should be noted that these trusted multi-view learning methods heavily rely on high-quality ground-truth labels, which may not always be available or reliable in real-world scenarios. This limitation motivates for us to delve into the problem of GTML, which aims to learn a reliable multi-view learning model under the guidance of noisy labels.
@@ 3.1 Notations and Problem Statement
We use $\{\bm{x}^{v}_{n}\in\mathbb{R}^{d_{v}}\}_{v=1}^{V}$ to denote the feature of the $n$-th instance, which contains $V$ views, $d_{v}$ denotes the dimension of the $v$-th view. $y_{n}\in\{1,...,K\}$ denotes the ground-truth category, where $K$ is the number of all categories. In the generalized trusted multi-view classification problem, the labels of some data instances contain noise as shown in Figure 2. Therefore, we use $\{\tilde{y}_{n}\in\{1,...,K\}\}_{n=1}^{N}$ as the noisy label set that may have been corrupted.
The objective is to learn a trusted classification model according to noisy training instances ${\{\{\bm{x}_{n}^{v}\}_{v=1}^{V}},\tilde{y}_{n}\}_{n=1}^{N_{train}}$. For the instances of the test sets, the model should predict the category $\{{y}_{n}\}$ and uncertainty $\{{u}_{n}\}$, which can quantify the uncertainty caused by low-quality feature and noisy label.
@@ 3.2.3 Trust Evidential Multi-view Fusion
After obtaining opinions from multiple views, we consider dynamically integrating them based on uncertainty to produce a combined opinion. We achieve this via the Dempster’s combination rule. We take the clean opinion aggregation as example. Given clean opinions of two views of the same instance, i.e., ${\mathcal{O}}^{1}=(\bm{b}^{1},u^{1})$ and ${\mathcal{O}}^{2}=(\bm{b}^{2},u^{2})$, the computation to obtain the aggregated opinion ${\mathcal{O}}=\mathcal{O}^{1}\diamond\mathcal{O}^{2}=(\bm{b},u)$ is defined as follows:
@@ 3.3.1 Classification loss
We capture view-specific evidence from a single view $\bm{x}^{v}$ of a sample. The vector $\bm{e}^{v}=f^{v}(\bm{x}^{v})$ denotes the clean class-posterior evidence obtained from the corresponding view network prediction. This evidence undergoes correction through Eq. (4) to yield the noisy class-posterior evidence, denoted as $\tilde{\bm{e}}^{v}$, along with the associated Dirichlet parameter $\tilde{\bm{\alpha}}^{v}$. Then in the constructed inference framework, the evidence multi-classification loss $\mathcal{L}$ containing the classification loss $\mathcal{L}_{ace}$ and the Kullback-Leibler (KL) divergence term $\mathcal{L}_{KL}$ is defined, where the classification loss $\mathcal{L}_{ace}$ is obtained by adjusting for the conventional cross-entropy loss, as denoted below:
@@ 3.3.2 Uncertainty-guided correlation loss
Non-diagonal elements. The constraints on the diagonal elements could be regarded as guiding the probability of the sample being mislabeled. The probability of being mislabeled as another category is influenced by the inherent relationship between the different categories. For example, “dog” is more likely to be labeled as “wolf” than as “plane”. Considering that samples in the same class can be easily labeled as the same error class, the transfer probabilities of their non-diagonal elements should be close. In addition, since the labeled information may contain noise, we aim to eliminate the misleading of the error samples in the same class. To solve this problem, we construct the affinity matrix $\{S\}_{v=1}^{V}$ for each viewpoint and calculate this loss with only the $k$ most similar samples in the same class, and for the $n$-th sample:
@@
$s_{mn}^{v}=\left\{\begin{array}[]{ cc}e^{-\frac{||\bm{x}_{m}^{v}-\bm{x}_{n}^{v}||^{2}}{\sigma^{2}}},&\text{if}\ \bm{x}_{m}^{v}\in\mathcal{N}(\bm{x}_{n}^{v},k)\ \text{and}\ \tilde{y}_{n}=\tilde{y}_{m},\\ 0,&\text{else},\end{array}\right.$ (11)
where $s_{nm}^{v}$ denotes the $(n,m)$-th element in the affinity matrix ${S}^{v}$ for the $v$-th view, which measures the similarity between $\bm{x}_{n}^{v}$ and $\bm{x}_{m}^{v}$. $\mathcal{N}(\bm{x}_{p}^{v},k)$ indicates the $k$-nearest neighbours of the view $\bm{x}_{p}^{v}$. $\bar{\bm{T}}^{v}_{i}$ is the matrix after zeroing the diagonal elements of $\bm{T}^{v}_{i}$. Thus, the overall regularization term for the inter-sample uncertainty bootstrap is expressed as:
@@ 4 Experiments
In this section, we test the effectiveness of the proposed method on 5 real-world multi-view datasets with different proportions of labelling noise added. In addition, we also verify the ability of the model to handle low-quality features and noisy label.

### v2

@@ Abstract
Multi-view learning methods often focus on improving decision accuracy while neglecting the decision uncertainty, which significantly restricts their applications in safety-critical applications. To address this issue, researchers propose trusted multi-view methods that learn the class distribution for each instance, enabling the estimation of classification probabilities and uncertainty. However, these methods heavily rely on high-quality ground-truth labels. This motivates us to delve into a new generalized trusted multi-view learning problem: how to develop a reliable multi-view learning model under the guidance of noisy labels? We propose a trusted multi-view noise refining method to solve this problem. We first construct view-opinions using evidential deep neural networks, which consist of belief mass vectors and uncertainty estimates. Subsequently, we design view-specific noise correlation matrices that transform the original opinions into noisy opinions aligned with the noisy labels. Considering label noises originating from low-quality data features and easily-confused classes, we ensure that the diagonal elements of these matrices are inversely proportional to the uncertainty, while incorporating class relations into the off-diagonal elements. Finally, we aggregate the noisy opinions and employ a generalized maximum likelihood loss on the aggregated opinion for model training, guided by the noisy labels. We empirically compare TMNR with state-of-the-art trusted multi-view learning and label noise learning baselines on 5 publicly available datasets. Experiment results show that TMNR outperforms baseline methods on accuracy, reliability and robustness. The code and appendix are released at https://github.com/YilinZhang107/TMNR.
@@ 1 Introduction
Multi-view data is widely present in various real-world scenarios. For instance, in the field of healthcare, a patient’s comprehensive condition can be reflected through multiple types of examinations; social media applications often include multi-modal contents such as textual and visual reviews Liu et al. (2024). Multi-view learning synthesizes both consistency and complementary information to obtain a more comprehensive understanding of the data. It has generated significant and wide-ranging influence across multiple research areas, including classification Chen et al. (2024); Wang et al. (2022), clustering Xu et al. (2023); Huang et al. (2023); Wen et al. (2022), recommendation systems Lin et al. (2023); Nikzad-Khasmakhi et al. (2021), and large language models Min et al. (2023).
@@
Figure 1: The generalized trusted multi-view learning problem: the model should recognize the feature and model uncertainty caused by low-quality feature and noisy labels, respectively.
@@
Most existing multi-view learning methods focus on improving decision accuracy while neglecting the decision uncertainty. This significantly limits the application of multi-view learning in safety-critical scenes, such as healthcare. Recently, Han et al. propose a pioneering work Han et al. (2020), Trusted Multi-view Classification (TMC), to solve this problem. TMC calculates and aggregates the evidences of all views from the original data features. It then utilizes these evidences to parameterize the class distribution, which could be used to estimate the class probabilities and uncertainty. To train the entire model, TMC requires the estimated class probabilities to be consistent with the ground-truth labels. Following this line, researchers propose novel evidence aggregation methods, aiming to enhance the reliability and robustness in the presence of feature noise Gan et al. (2021); Qin et al. (2022), conflictive views Xu et al. (2024) and incomplete views Xie et al. (2023).
@@
Regretfully, these trusted multi-view learning methods consistently rely on high-quality ground-truth labels. The labeling task is time-consuming and expensive especially when dealing with large scale datasets, such as user generated multi-modal contents in social media applications. This motivates us to delve into a new Generalized Trusted Multi-view Learning (GTML) problem: how to develop a reliable multi-view learning model under the guidance of noisy labels? This problem encompasses two key objectives: 1) detecting and refining the noisy labels during the training stage; 2) recognizing the model’s uncertainty caused by noisy labels. For example, instances belonging to classes like “dog” and “wolf” might exhibit similarities and are prone to being mislabelled. Consequently, the model should exhibit higher decision uncertainty in such cases. An intuitive analogy is an intern animal researcher (model) may not make high-confidence decisions for all animals (instances), but is aware of the cases where a definitive decision is challenging.
@@
In this paper, we propose an Trusted Multi-view Noise Refining (TMNR) method for the GTML problem. We consider the label noises arising from two sources: low-quality data features, such as blurred and incomplete features, and easily-confused classes, such as classes “dog” and “wolf”. Our objective is to leverage multi-view consistent information for noise detection. To achieve this, we first construct the view-specific evidential Deep Neural Networks (DNNs) to learn view-specific evidence, which could be termed as the amount of support to each category collected from data. We then model the view-specific distributions of class probabilities using the Dirichlet distribution, parameterized with view-specific evidence. These distributions allow us to construct opinions, which consist of belief mass vectors and uncertainty estimates. We design view-specific noise correlation matrices to transform the original opinions to the noisy opinions, which aligns with the noisy labels. Considering that low-quality data features are prone to mislabeling, we require the diagonal elements of the noise correlation matrices to be inversely proportional to the uncertainty. Additionally, we incorporate class relations into the off-diagonal elements. For instance, the elements corresponding to “dog” and “wolf” should have larger values since these two classes are easily mislabelled. Next, we aggregate the noisy opinions to obtain the common evidence. Finally, we employ a generalized maximum likelihood loss on the common evidence, guided by the noisy labels, for model training.
@@ 2.1 Deep Multi-view Fusion
Multi-view fusion has demonstrated its superior performance in various tasks by effectively combining information from multiple sources or modalities Liang et al. (2021); Zhou et al. (2023). According to the fusion strategy, existing deep multi-view fusion methods can be roughly classified into feature fusion Hu et al. (2024); Xu et al. (2022); Liu et al. (2023) and decision fusion Jillani et al. (2020); Liu et al. (2022). A major challenge of feature fusion methods is each view might exhibit different types and different levels of noise at different points in time. Trust decision fusion methods solve this making view-specific trust decisions to obtain the view-specific reliabilities, then assigning large weights to these views with high reliability in the multi-view fusion stage. Following this line, Xie et al. Xie et al. (2023) tackle the challenge of incomplete multi-view classification through a two-stage approach, involving completion and evidential fusion. Xu et al. Xu et al. (2024) focus on making trust decisions for instances that exhibit conflicting information across multiple views. They propose an effective strategy for aggregating conflicting opinions and theoretically prove this strategy can exactly model the relation of multi-view common and view-specific reliabilities. However, it should be noted that these trusted multi-view learning methods heavily rely on high-quality ground-truth labels, which may not always be available or reliable in real-world scenarios. This limitation motivates for us to delve into the problem of GTML, which aims to learn a reliable multi-view learning model under the guidance of noisy labels.
@@ 3.1 Notations and Problem Statement
We use $\{\bm{x}^{v}_{n}\in\mathbb{R}^{d_{v}}\}_{v=1}^{V}$ to denote the feature of the $n$-th instance, which contains $V$ views, $d_{v}$ denotes the dimension of the $v$-th view. $y_{n}\in\{1,...,K\}$ denotes the ground-truth category, where $K$ is the number of all categories. In the generalized trusted multi-view classification problem, the labels of some data instances contain noise as shown in Figure 2. Therefore, we utilize $\{\tilde{y}_{n}\in\{1,...,K\}\}_{n=1}^{N}$ as the set of noisy labels that may have been corrupted
The objective is to learn a trusted classification model according to noisy training instances ${\{\{\bm{x}_{n}^{v}\}_{v=1}^{V}},\tilde{y}_{n}\}_{n=1}^{N_{train}}$. For the instances of the test sets, the model should predict the category $\{{y}_{n}\}$ and uncertainty $\{{u}_{n}\}$, which can quantify the uncertainty caused by low-quality feature and noisy labels.
@@ 3.2.3 Trust Evidential Multi-view Fusion
After obtaining opinions from multiple views, we consider dynamically integrating them based on uncertainty to produce a combined opinion. We achieve this via the Dempster’s combination rule Jøsang (2016). We take the clean opinion aggregation as example. Given clean opinions of two views of the same instance, i.e., ${\mathcal{O}}^{1}=(\bm{b}^{1},u^{1})$ and ${\mathcal{O}}^{2}=(\bm{b}^{2},u^{2})$, the computation to obtain the aggregated opinion ${\mathcal{O}}=\mathcal{O}^{1}\diamond\mathcal{O}^{2}=(\bm{b},u)$ is defined as follows:
@@ 3.3.1 Classification loss
We capture view-specific evidence from a single view $\bm{x}^{v}$ of a sample. The vector $\bm{e}^{v}=f^{v}(\bm{x}^{v})$ denotes the clean class-posterior evidence obtained from the corresponding view network prediction. This evidence undergoes correction through Eq. (4) to yield the noisy class-posterior evidence, denoted as $\tilde{\bm{e}}^{v}$, along with the associated Dirichlet parameter $\tilde{\bm{\alpha}}^{v}$. Then in the constructed inference framework, the evidence multi-classification loss $\mathcal{L}$ containing the classification loss $\mathcal{L}_{ace}$ and the Kullback-Leibler (KL) divergence term $\mathcal{L}_{KL}$ is defined, where the classification loss $\mathcal{L}_{ace}$ is obtained by adjusting for the conventional cross-entropy loss, i.e., the generalized maximum likelihood loss, as follows:
@@ 3.3.2 Uncertainty-guided correlation loss
Non-diagonal elements. The constraints on the diagonal elements could be regarded as guiding the probability of the sample being mislabeled. The probability of being mislabeled as another category is influenced by the inherent relationship between the different categories. For example, “dog” is more likely to be labeled as “wolf” than as “plane”. Considering that samples in the same class can be easily labeled as the same error class, the transfer probabilities of their non-diagonal elements should be close. In addition, since the labeled information may contain noise, we aim to eliminate the misleading of the error samples in the same class. To solve this problem, we construct the affinity matrix $\{\bm{S}^{v}\}_{v=1}^{V}$ for each viewpoint and calculate this loss with only the $k$ most similar samples in the same class, and for the $n$-th sample:
@@
$s_{nm}^{v}=\left\{\begin{array}[]{ cc}e^{-\frac{||\bm{x}_{n}^{v}-\bm{x}_{m}^{v}||^{2}}{\sigma^{2}}},&\text{if}\ \bm{x}_{m}^{v}\in\mathcal{N}(\bm{x}_{n}^{v},k)\ \text{and}\ \tilde{y}_{n}=\tilde{y}_{m},\\ 0,&\text{else},\end{array}\right.$ (11)
where $s_{nm}^{v}$ denotes the $(n,m)$-th element in the affinity matrix ${\bm{S}}^{v}$ for the $v$-th view, which measures the similarity between $\bm{x}_{n}^{v}$ and $\bm{x}_{m}^{v}$. $\mathcal{N}(\bm{x}_{n}^{v},k)$ indicates the $k$-nearest neighbours of the view $\bm{x}_{n}^{v}$. $\bar{\bm{T}}^{v}_{i}$ is the matrix after zeroing the diagonal elements of $\bm{T}^{v}_{i}$. Thus, the overall regularization term for the inter-sample uncertainty bootstrap is expressed as:
@@ 4 Experiments
In this section, we test the effectiveness of the proposed method on 5 real-world multi-view datasets with different proportions of labelling noise added. In addition, we also verify the ability of the model to handle low-quality features and noisy labels.
@@ 5 Conclusion
Acknowledgments
This research was supported by the National Natural Science Foundation of China (Grant Nos. 62133012, 61936006, 62103314, 62073255, 62303366), the Key Research and Development Program of Shanxi (Program No. 2020ZDLGY04-07), Innovation Capability Support Program of Shanxi (Program No. 2021TD-05) and Natural Science Basic Research Program of Shaanxi under Grant No.2023-JC-QN-0648.
Appendix
In this appendix, we provide a more detailed description of the noise correlation matrix and types of label-noise, and show more implementation details of the proposed Trusted Multi-view Noise Refining (TMNR) method, hyper-parameter settings and performance comparisons under different situations.
A    Label-Noise Learning
A.1    Types of Label-Noise
In Class-Conditional Noise (CCN), the process of label corruption is independent of the features of the instance itself. Its ground-truth labels are corrupted by a dataset-specific noise correlation matrix $\bm{T}=[t_{ij}]_{i,j=1}^{K}\in[0,1]^{K\times K}$, where $t_{ij}=P(\tilde{y}=j|y=i)$ , $i,j\in\{1,...,K\}$ and $K$ is the number of all categories. In this hypothesis, when the noise rate is $\sigma$, the noise is said to be symmetric noise (as in Figure 6(a)) if $\forall_{i=j}\ t_{ij}=(1-\sigma)$ and $\forall_{i\neq j}\ t_{ij}=(\sigma/(K-1))$. As opposed to this, asymmetric noise means that for $\forall_{i}$, ${\textstyle\sum_{j=1,j\neq i}^{K}}t_{ij}=\sigma$ and $\exists_{j\neq i}\ t_{ij}\neq[\sigma/(K-1)]$, then it means that when labels are incorrectly labelled, it is more likely that they will be corrupted to a specific category rather than randomly labelled into other categories with average probability. When a category of labels will only transform into a specific category of labels, as in $\forall_{i=j}\ t_{ij}=(1-\sigma)$ and $\exists_{j\neq i}t_{ij}=\sigma$, then it is called flip noise, shown in Figure 6(b).
In the real world, considering that samples with ambiguous features or poor collection quality are more likely to be mislabeled. Therefore, in Instance-Dependent Noise (IDN) modeling, it is assumed that the process of labeling corruption depends on its own features and category labels, and the sample $\bm{x}_{n}$ specific noise correlation matrix is defined as $t_{ij}=P(\tilde{y}=j|y=i,\bm{x}_{n})$.
(a) Symmetric Noise
(b) Flip Noise
Figure 6: Different types of CCN correlation matrices when the noise ratio $\sigma=20\%$ and $K=3$.
B    Supplementary Experimental Content
B.1    Noise Generation
To destroy the labels of the clean dataset used in the main text, we emulate the previous method Cheng et al. [2020] of destroying labels to generate datasets containing instance-dependent label noise. First we train a trusted classifier $g(\cdot)$ on a subset of the clean dataset using the same model as the Evidence Extraction Network part and make predictions for each sample that will be used for training. Then the noisy labels are obtained by taking into account the amount of evidence $\bm{e}$ obtained for each class and the uncertainty ${u}$ of the opinions. Specifically, we measure the magnitude of the probability of whether a sample is corrupted in terms of uncertainty, i.e., elements with greater uncertainty are more likely to be corrupted. When a sample is selected for destruction, we pick $\tilde{y}=arg\ max(e_{k})$ and $k\neq y$, $\ k\in\{1,...,K\}$. This means that we set the category that the classifier believes to be most similar to the category to which the instance belongs as the noise label to satisfy our assumption of instance-related label noise. We added noise with a scale of 10% to 50% to all the datasets to evaluate the proposed method.
Datasets Size Classes Dimensionality
UCI 2000 10 6/47/24
PIE 680 68 484/256/279
BBC 685 5 4659/4633/4665/4684
Caltech101 8677 101 48/40/254/1984/512/928
Leaves100 1600 100 64/64/64
Table 3: Summary of feature dimensions for each view for all datasets.
Datasets Method Instance-Dependent Noise
[…]
PIE - - 89.85$\pm$2.69 84.41$\pm$3.13 82.65$\pm$3.24 77.65$\pm$2.11 66.47$\pm$3.50 60.44$\pm$5.82
- 89.26$\pm$2.05 85.00$\pm$2.61 82.79$\pm$3.04 77.65$\pm$1.51 67.79$\pm$3.63 60.00$\pm$6.19
- 89.71$\pm$2.23 84.41$\pm$3.06 82.79$\pm$3.28 77.06$\pm$1.43 67.94$\pm$3.44 60.88$\pm$4.25
90.74$\pm$1.51 85.51$\pm$1.95 82.94$\pm$2.92 78.38$\pm$1.10 68.68$\pm$3.71 61.47$\pm$4.82
Caltech101 - - 90.38$\pm$0.87 90.75$\pm$0.55 86.86$\pm$0.95 87.57$\pm$1.79 79.41$\pm$1.11 70.67$\pm$2.68
- 90.79$\pm$0.79 90.59$\pm$1.08 87.82$\pm$1.29 87.62$\pm$1.13 79.62$\pm$1.33 71.21$\pm$2.35
- 90.21$\pm$1.00 90.75$\pm$0.79 87.62$\pm$0.73 88.03$\pm$0.86 80.46$\pm$1.94 73.64$\pm$2.87
90.63$\pm$0.83 91.05$\pm$0.83 88.08$\pm$0.98 88.08$\pm$1.15 80.63$\pm$1.74 73.93$\pm$3.00
Leaves100 - - 63.75$\pm$2.70 61.19$\pm$1.50 61.56$\pm$1.98 55.50$\pm$3.30 55.31$\pm$6.24 47.56$\pm$6.62
- 69.25$\pm$3.62 66.00$\pm$2.94 63.19$\pm$2.20 58.06$\pm$5.90 56.25$\pm$5.94 52.56$\pm$5.64
- 68.56$\pm$2.76 70.50$\pm$3.66 64.62$\pm$2.25 61.69$\pm$4.23 58.06$\pm$2.48 58.81$\pm$4.81
69.19$\pm$2.15 70.81$\pm$3.17 68.25$\pm$2.23 65.00$\pm$4.53 61.81$\pm$4.71 59.50$\pm$3.17
Table 4: Ablation study on three datasets. $\mathcal{L}_{acc}$ denotes overall classification loss, $\mathcal{M}$ indicates uncertainty bootstrap regularization and $\mathcal{L}_{con}$ denotes loss of consistency of the inter-view correlation matrix. “✔” indicates that the corresponding component in the TMNR is applied, and “-” indicates that it is not used. Best results are highlighted by bold.
B.2    Network Construction
The UCI Dataset, PIE Dataset, BBC Dataset, Caltech101 Dataset and Leaves100 Dataset are composed of pre-extracted vectorized features. Then we extract view-specific evidence using a fully connected layer network with ReLU activation functions. To be fair, for each dataset we add only one fully connected layer for each view with inputs as the feature dimensions of the view and outputs as the number of categories $K$. The feature dimensions of each view for all datasets are summarised in Table 3. For all datasets, the hyper-parameter $\beta$ is set to $1e^{-2}$, and the non-diagonal elements of the correlation matrix are compared with their $5$-nearest neighbors in the computation of the uncertainty bootstrap loss $\mathcal{M}$. The selection process of hyper-parameters $\gamma$ for different datasets will be shown later.
B.3    Ablation Study
(a) BBC IDN-$10\%$
(b) BBC IDN-$40\%$
Figure 7: Evaluating the performance of TMNR with its different degradation methods in terms of uncertainty assessment on BBC datasets containing 10% and 40% noise.
Figure 8: Classification accuracy when adjusting $\gamma$ on all datasets with 30% noise rate.
(a) UCI
(b) PIE
(c) Leaves100
(d) Caltech101
Figure 9: Identification of in/out-of-distribution samples.
To demonstrate the effectiveness of our proposed uncertainty bootstrap regularisation $\mathcal{M}$ and loss of consistency of inter-view correlation matrices $\mathcal{L}_{con}$ in optimising noisy correlation matrices. We performed an ablation study: without adding any additional assumptions to the correlation matrix and using only one of the two modules. We verify their effectiveness by evaluating their performance under different levels of labelled noise. As shown in Table 4, each constraint almost always improves the performance of the experimental results under different noise levels, illustrating that the correlation patterns between clean and noisy evidence distributions can be learnt more effectively through uncertainty bootstrapping and inter-perspective consistency constraints. And the constraints on the correlation matrix work better when the noise percentage is high.
Meanwhile, in order to verify the effectiveness of the proposed noise correlation matrix in eliminating the effect of label noise, we obtain TMNR-NG and TMNR-NT by removing the additional constraints and correlation matrices sequentially. The uncertainty distributions obtained from the prediction of the test set using the TMNR with the two degradation methods trained on the BBC dataset containing 10% and 40% label noise are shown in Figure 7. The results show that although the uncertainty of the model inevitably increases with the increase of the noise ratio in the training data, the incorporation of the correlation matrix effectively reduces the uncertainty of the test sample decisions, especially when applying the constraints we designed. Lower uncertainty implies sharper Dirichlet distributions corresponding to the aggregation of opinions from multiple perspectives, indicating that sufficient evidence has been observed to make trusted decisions.
B.4    Hyper-parametric Analysis
We analyze it here the sensitivity of the hyper-parameter $\gamma$ adjusting for the inter-view correlation matrix consistency loss $\mathcal{L}_{con}$ on all datasets containing 30% noise. The results is shown in Figure 8. We observe that although different datasets have different sensitivities to the parameter $\gamma$, appropriate parameter values can still improve the overall performance of the model. We set different $\gamma$ values for each dataset based on the results of this experiment. The UCI dataset, PIE dataset, BBC dataset and the Leaves100 dataset are set to $1e^{4}$, and the Caltech101 dataset is set to $1e^{3}$.
B.5    Identification of Out-of-Distribution Data
To verify the effectiveness of our proposed TMNR as a trusted model in data noise identification, we add Gaussian noise with fixed standard deviation ($\sigma=10$) to $50\%$ of the test samples in the four datasets so that they constitute out-of-distribution (OOD) samples, and the remaining data serve as in-distribution samples. Their uncertainties were predicted using the model obtained on the training data without labeling noise and the results are shown in Figure 9. We can observe that the intra-distribution samples all obtained lower uncertainty than the OOD samples on all datasets. Meanwhile, this UCI dataset with higher prediction accuracy exhibits lower uncertainty overall, while the Leaves100 dataset with lower prediction accuracy has higher uncertainty. These results prove the reasonableness and effectiveness of TMNR’s measure of data uncertainty to ensure that the decisions output by our model are trustworthy.